Present: Rich Bennett, Will Chaney, Bill Covey, Debra Harris, Winston Harris, Bobbie Meng, Everett Phillips, Suzy Shaw, Hilton Soto, Ying TangAnnouncements
Ongoing Systems Updates
- Bobbie received an email from a former Systems employee, Paul Fortier, announcing that he has recently become a grandfather to his son's new baby girl.
- 60 Cow order to go out today (replacements for 75/100/SNA).
- 6 more hard drives to be ordered today.
- Distribution list from Martha for the 20 Cows received last week has been issued. Need to verify with Leilani that they have space to store if not install the new computers they are scheduled to receive.
- The Control Unit currently parked in front of the inside vault needs to be surveyed. The space will be needed upon arrival of the 60 Cows.
Software Projects Status
- NT Security Packs - go to 6a and add the patch. Should be up on NT2 by this afternoon in the Systems Folder.
- Virus software - need to keep updated. There are some new viruses out there that will attack your email address book. There have been some hacks on campus mail lists. We are running McAfee net scan on the servers.
- There was some instability with the Smathers Novell server on Friday night. Tom Caswell had to bring it back up. There was no apparent reason for the instability, but hopefully this type of occurrence will diminish when we bring it up to Y2K compliance soon.
- Smathers network reconfiguration - will hold off until January.
- NT 5 replacement - Will is working on it, had some problems on Friday when it died on him and he had to rebuild it. He's looking for an automatic way to transfer files over and is looking to have it switched out by the end of the week.
- Wiring - Will is scheduled to meet with contractors tomorrow at 10 AM to discuss estimates on the Documents wiring job. Mention was also made about rearranging the West wiring closet.
- Wiring database - Ying is making some enhancements on it - should be ready to demo within the next day or two.
- Forms to mail - Suzy got the port to Linux working at home and NERDC, we will run the same thing here. Will need send mail & APACHE web server. Rich & Suzy will be changing all the URLs to a generic address. Will be changed in the database - no recoding.
- NERDC/OCLC tape conversion - Sharon Wheeler is working on it at NERDC, she is running into problems with very long and some bad files. The project is ongoing.
- Q&A database - need to move to current projects, still working on requirements. Winston contacted Don Frick about providing item records to coordinate with the Q&A database.
- ATC Standalone - ready for multi-user testing. Winston set up the system to download, then email him when it's finished. It was scheduled for 10PM last night, but was still running at 10AM this morning. It was only 1/5th finished after 12 hours. FCLA should be moving WEBLUIS to a new machine.
- Hilton has modified the Goza database to SQL server from Caroline. There is only one other old newspaper database on Caroline. Most of the info on it is located somewhere else. We will save the database, but won't convert it unless it becomes necessary.
- Hilton has the web-based Trouble Ticket Viewer almost complete. He is using InterDev.
- Compustat tapes - haven't heard from Peter McKay regarding the beta-testing of the CDs. Debra will follow-up.
- Dead Week on Thursday & Friday, finals next week. Should have student coverage through Wednesday of next week, then they'll start disappearing.
- Students don't return until January 10 - they have an extra week off this year - we'll be without them for approximately a month.
- Suzy out December 9 & 10.
- Library Christmas Party - Wednesday, Dec. 15 (2:00-4:00 PM).
- Rich out December 20-29.
- Christmas Holiday December 24.
- Will out December 28-29.
- New Year Holiday - Friday, Jan. 31.
- Everett out January 6-12.
- Everett out January 28-31.
Back to the Systems Homepage
Last updated December 13, 1999
by Debra Harris